A $25,000 10% bond redeemable at par on December 1, 2028, is purchased on September 25, 2017, to yield 7.6% compounded semi-annually. Bond interest is payable semi-annually. Find answer using financial calculator.

a) What is the cash price of the bond?

b) What is the accrued interest?

c) What is the quoted price?
Answer #1

a: Using financial calculator
Input: FV= 25000, N=23

I/Y=7.6/2= 3.8

PMT=10%*25000/2 = 1250

Solve for PV as -29546.64

Cash price = $29546.64

b: Accrued interest = from 1 May to 25 Sep. (including both days)

=10%*25000*148/365

= 1013.70

c: Quoted price= 29546.64+1013.70 = $30560.33